OpenStudy (anonymous):

Please help! Make a scenerio for me. Its 11 o'clock, im tired, & i haven't even eaten yet! Pleaseeeeeee! I need it soon too because i cant exit out of the test either or i will have to do all 20+ questions all over again. Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

jimthompson5910 (jim_thompson5910):

let's say you have pencils and pens

jimthompson5910 (jim_thompson5910):

let x = cost per pencil y = cost per pen if you have 2 pencils and 1 pen, and the total cost is $10, then 2x + y = 10 if you have 3 pencils and 4 pens, and the total cost is $25, then 3x + 4y = 25

jimthompson5910 (jim_thompson5910):

so that would give you the system 2x + y = 10 3x + 4y = 25

jimthompson5910 (jim_thompson5910):

let's solve for x and y

jimthompson5910 (jim_thompson5910):

2x + y = 10 y = 10 -2x ---------------- 3x + 4y = 25 3x + 4(10-2x) = 25 3x + 40 - 8x = 25 -5x+40=25 -5x=25-40 -5x=-15 x=(-15)/(-5) x=3 y = 10 - 2x y = 10 - 2(3) y = 10 - 6 y = 4 So each pencil costs $3 and each pen costs $4 each

OpenStudy (anonymous):

thank you sooooo much. I can go to bed after one more question. Can ya help?

jimthompson5910 (jim_thompson5910):

sure just one more though

OpenStudy (anonymous):

promise!

OpenStudy (anonymous):

You graph a system of equations and the two lines intersect one another. Explain what this tells you about the: solution to the system the two equations the result if the system was solved algebraically

jimthompson5910 (jim_thompson5910):

go for it

jimthompson5910 (jim_thompson5910):

there is exactly one solution this is because a solution corresponds to where the two lines cross

jimthompson5910 (jim_thompson5910):

they cross at exactly one point, so there is exactly one solution

jimthompson5910 (jim_thompson5910):

this means that the system of equations are independent and consistent

jimthompson5910 (jim_thompson5910):

was that all you needed?

OpenStudy (anonymous):

uhm could you rephrase that as 1. ( ) 2. ( ) 3. ( )

jimthompson5910 (jim_thompson5910):

1. exactly one solution 2. the system of equations are independent and consistent 3. one result because there is only one solution

jimthompson5910 (jim_thompson5910):

the reasoning is explained above

OpenStudy (anonymous):

thanks so much. Thats all i need. You were so helpful, im definitely becoming a fan. Now i can eat and go to bed. Goodnight and once again thank you sooo much

jimthompson5910 (jim_thompson5910):

good night, yw